# GenOM: Ontology Matching with Description Generation and Large Language Model

### v2

@@ Abstract
Ontology matching (OM) plays an essential role in enabling semantic interoperability and integration across heterogeneous knowledge sources, particularly in the biomedical domain which contains numerous complex concepts related to diseases and pharmaceuticals. This paper introduces GenOM, a large language model (LLM)-based ontology alignment framework, which enriches the semantic representations of ontology concepts via generating textual definitions, retrieves alignment candidates with an embedding model, and incorporates exact matching-based tools to improve precision. Extensive experiments conducted on the OAEI Bio-ML track demonstrate that GenOM can often achieve competitive performance, surpassing many baselines including traditional OM systems and recent LLM-based methods. Further ablation studies confirm the effectiveness of semantic enrichment and few-shot prompting, highlighting the framework’s robustness and adaptability.
@@
1{jiaoyan.chen, Renate.Schmidt}@manchester.ac.uk
1yiping.song@postgrad.manchester.ac.uk
@@ 1 Introduction
In recent years, the rapid growth of domain-specific ontologies has intensified the demand for semantic interoperability and knowledge integration across heterogeneous information systems [1]. In particular, the biomedical domain [2] has witnessed the proliferation of large and complex ontologies, such as those used to model diseases, anatomy, and clinical procedures, which further amplifies this challenge. While ontologies serve as structured and machine-interpretable representations of conceptual knowledge, they are often engineered independently by different research communities or organisations. Each ontology typically reflects particular domain-specific objectives, modelling assumptions, and terminological conventions, which collectively result in substantial structural and semantic heterogeneity. This diversity, though valuable for capturing local perspectives, poses significant challenges to semantic interoperability and the effective integration of knowledge across heterogeneous information systems. Consequently, bridging conceptual discrepancies among ontologies has become a central challenge for achieving seamless knowledge exchange and reuse in data-intensive environments. Against this backdrop, the goal of this paper is to develop an ontology matching (OM) framework that can effectively bridge heterogeneous ontologies by leveraging the reasoning and generation capabilities of large language models (LLMs).
@@
As a core mechanism for achieving this goal, ontology matching (OM) has emerged as a pivotal research area in response to the growing need for semantic integration across heterogeneous knowledge bases [3, 4]. Its primary objective is to uncover semantic correspondences between entities that originate from different ontologies, thereby enabling the consistent interpretation and exchange of knowledge. Through establishing formal relations such as equivalence, where two entities express the same or highly similar meanings, and subsumption, where one entity denotes a more general or more specific concept, OM serves as a cornerstone for semantic interoperability in distributed information environments.
@@
Nevertheless, achieving accurate alignments remains highly challenging due to the intrinsic heterogeneity of ontologies. This heterogeneity is manifested in several forms. First, terminological variation arises when identical concepts are described using distinct lexical labels or synonyms. Second, structural variation reflects the diversity of ontology design choices, ranging from deep hierarchical taxonomies to flat enumerative structures. Third, granularity variation occurs when similar knowledge is modelled at differing levels of abstraction. Together, these inconsistencies amplify both the cognitive and computational complexity of the matching task. The challenge becomes even more pronounced as ontologies continue to scale. For example, SNOMED CT [5], one of the largest biomedical terminologies, comprises hundreds of thousands of concepts, which has led some studies to further decompose it into smaller, task-specific subontologies for improved tractability [6, 7], yet purely manual alignment remains impractical at this scale, further motivating the need for scalable automated solutions.
@@
Over the past decade, a range of automated systems, such as LogMap [8] and AgreementMakerLight (AML) [9], have been proposed to address this challenge. These approaches integrate lexical similarity, structural heuristics, and background knowledge resources to identify candidate mappings. While effective in well-structured and lexically rich scenarios, they often fail to capture the deeper semantics encoded in complex concept descriptions. Recent progress in LLMs has fundamentally reshaped this landscape. By virtue of their powerful contextual reasoning and linguistic generalisation abilities, LLMs offer an unprecedented opportunity to represent and compare ontological knowledge at a semantic level. Several contemporary OM systems have begun to incorporate LLMs at different stages of the alignment pipeline, including semantic embedding and direct alignment prediction. For instance, LLM4OM [10] integrates ChatGPT with OpenAI embeddings to support pairwise concept matching, while Olala [11] leverages LLM-based semantic reasoning to assess concept equivalence. A more detailed review is given in Section 2. Despite these advances, current LLM-based approaches remain constrained by practical and methodological limitations. Some rely on extremely large models (e.g., the 70B-parameter LLaMA-2 employed in Olala [11]), which incur considerable computational costs and hinder scalability, whereas others [12] struggle to maintain robustness when confronted with complex or large-scale alignment scenarios.
@@
To address these limitations, we develop GenOM, an ontology matching framework that leverages LLMs for semantic enrichment, candidate generation, and mapping judgement. [^1] A preliminary version of this idea, focusing on LLM-based definition generation, was presented at the OM Workshop @ ISWC 2025 [13]. GenOM begins by extracting lexical and structural information from both source and target ontologies—including labels, synonyms, parent relations, and equivalence axioms—and employs an LLM-driven description generation module to produce semantically enriched textual definitions for each concept. These enriched descriptions are then embedded into a vector space to support efficient retrieval of semantically similar candidates, while an LLM-based mapping judgement component and a lightweight exact-matching module jointly refine the alignments.
@@
In this work, we instantiate GenOM using a diverse suite of LLMs, including Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Llama3-8B-Instruct, which vary in architectural design and parameter scale. This enables a systematic analysis of how model capacity influences definition quality, semantic representation, and overall alignment performance.
@@
Contributions. This work makes the following key contributions:
  • GenOM: an LLM-enhanced ontology matching framework. We introduce GenOM, a new ontology matching framework that integrates LLM-driven semantic enrichment with embedding-based candidate generation and LLM-based mapping judgement. GenOM achieves strong performance across multiple biomedical alignment tasks, demonstrating the effectiveness of combining LLM semantics with traditional matching components.
  • A novel evaluation methodology for assessing the quality of LLM-generated concept definitions. We propose a dedicated evaluation scheme for concept definitions in OM tasks, combining lexical metrics, alignment-based measures, and an LLM-as-a-judge semantic assessment to quantify the correctness and discriminativeness of generated definitions.
@@ 2 Related Work
At present, OM approaches can be broadly categorised into four main types: traditional knowledge-based systems, machine learning-based systems, pre-trained language model-based systems and more recently, large language models (LLMs)-based systems. Traditional systems, such as LogMap and AML, rely primarily on lexical similarity, structural heuristics, and external resources like UMLS or WordNet [14]. LogMap [8] extracts class names and searches for matches via external lexicons while addressing logical inconsistencies by selecting alignments with higher confidence scores. AML [9] employs multiple matching strategies, including exact and character-based matchers, and has shown strong performance on medical datasets. However, traditional ontology matching systems largely rely on string-based similarity measures and curated external lexicons, which have been shown to limit their coverage when semantically equivalent concepts differ substantially at the lexical level [15]. As a result, such approaches may fail to identify valid correspondences that are not explicitly recorded in external dictionaries or lexicons.
@@
With the rise of machine learning, several approaches have emerged that regard alignment as a classification or (embedding-based) similarity learning task [16, 17, 18, 19, 20, 21]. For example, DeepAlignment [17] vectorises class names and computes Euclidean distances to assess similarity, while the CNN-based system [21] uses character-level embeddings and hierarchical context to train binary classifiers for equivalence detection.
@@
In parallel, encoder-based pre-trained language models such as BERT [22] have been explored for ontology matching, leveraging contextualised representations learned from large-scale corpora. BioSTransformers [23] adopt a Siamese architecture based on domain-specific BERT models to compute semantic similarity between biomedical concepts. BERTMap [15] fine-tunes a domain-specific BERT model on ontology alignment corpora mainly automatically extracted from synonyms, enabling it to capture subtle semantic differences even when lexical overlap is low. This approach has demonstrated strong results, particularly in biomedical applications. BERTSubs [24] takes a similar architecture as BERTMap but focuses on the subsumption relationship.
@@
More recently, the emergence of generative LLMs, including the GPT series, the Llama series, and the T5 series, has stimulated a new line of research in ontology alignment. These LLMs provide strong contextual reasoning and linguistic generalisation capabilities, making them attractive tools for capturing semantic relations that go beyond lexical similarity. A number of studies have explored prompting strategies and retrieval-augmented pipelines to exploit these capacities. For example, Norouz et al. [12] employed GPT-4o in a prompt-driven alignment workflow, reporting strong recall, while highlighting challenges in reliably distinguishing equivalence from hierarchical relations. Yuan et al. [25] evaluated both open- and closed-source LLMs on biomedical alignment tasks, incorporating structural cues to strengthen predictive reliability. Hybrid frameworks that combine embedding-based similarity retrieval with LLM verification have also been proposed [10, 26], aiming to mitigate hallucination and improve alignment robustness. The Olala system [11] further integrates embedding-based filtering and post-processing modules with Llama-2, which is responsible for generating final alignment decisions. Hu et al. [27] have explored enriching ontological information by leveraging LLMs to expand or clarify incomplete concept metadata. Their approach employs LLaMA-3.3-70B-Instruct to, for instance, transform abbreviated forms into their likely full expressions and to generate explanatory descriptions for concepts with limited textual context.
Traditional ontology matching systems have also begun to incorporate LLMs to enhance their existing pipelines. For instance, the latest version of LogMap [28] integrates GPT-4-mini and models from the Gemini family to assist in resolving uncertain and ambiguous candidate mappings. This hybrid integration enables the system to refine difficult alignment cases that conventional lexical- or structure-based heuristics struggle to disambiguate, resulting in modest improvements over the original LogMap pipeline.
@@
Despite recent advances, existing LLM-based and LLM-augmented ontology matching approaches continue to face notable limitations. A central challenge arises from the fragmented and often sparse lexical information available within most ontologies [11, 10, 12]. LLMs must rely heavily on labels, short synonyms, and minimally structural cues, which provide insufficient semantic context for distinguishing between closely related or hierarchically similar medical concepts. As a result, current LLM-based systems often struggle in fine-grained disambiguation scenarios [10], where subtle terminological or relational differences must be accurately interpreted.
@@ 3.1 Task Formulation
OM is the process of discovering semantic correspondences between entities that belong to different ontologies. Formally, given two ontologies, the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$, let $C_{\text{s}}$ and $C_{\text{t}}$ denote sets of named concepts respectively contained in $O_{\text{s}}$ and $O_{\text{t}}$. The goal of OM is to identify a set of mappings that link semantically related concepts across the two ontologies. Each mapping consists of a concept pair $(c_{\text{s}},c_{\text{t}})$, where $c_{\text{s}}\in C_{\text{s}}$ and $c_{\text{t}}\in C_{\text{t}}$, together with an associated confidence value $\alpha_{st}\in[0,1]$, which quantifies the strength or likelihood of their semantic equivalence. The resulting alignment can therefore be expressed as a set of scored correspondences:
$M=\left\{(c_{\text{s}},c_{\text{t}},\alpha_{st})\mid c_{\text{s}}\in C_{\text{s}},\,c_{\text{t}}\in C_{\text{t}},\,\alpha_{st}\in[0,1]\right\}$ (1)
A higher value of $\alpha_{st}$ indicates greater semantic proximity between $c_{\text{s}}$ and $c_{\text{t}}$, and is often used as a confidence score for ranking candidate mappings or filtering high-confidence alignments. In practical settings, the OM task can encompass various types of semantic relations, such as equivalence, subsumption, and relatedness, depending on the intended application or evaluation protocol. Accordingly, the OM task can be regarded as a process of semantic reconciliation that transforms heterogeneous conceptual representations into a shared and interoperable knowledge space.
@@ 3.2 The Workflow of the GenOM Framework
The overall workflow of GenOM, illustrated in Figure 1, comprises five main components. 1) Ontology data extraction collects structural and lexical information for each named concept in the source and target ontologies, including labels, synonyms, parent concepts, and equivalence axioms. 2) An LLM-based definition generation module produces natural language definitions based on paraphrased descriptions of concepts using the extracted information. This enriches their semantic representations—particularly for entities without explicit textual descriptions [29, 30]. 3) A candidate generation stage employs an embedding-based retrieval process to compute pairwise cosine similarities between source and target concepts, thereby identifying the top-$k$ candidate mappings. 4) An LLM-based equivalence judgement module evaluates each candidate pair, determining whether the two concepts convey equivalent meanings by leveraging their generated definitions and structural context. 5) A post-processing and result fusion stage integrates the LLM prediction scores with cosine similarities and exact lexical matching results, retaining only high-confidence correspondences.
@@
Both lexical and structural characteristics of each concept from the source and target ontologies are extracted and utilized to support the following alignment process. Specifically, the extracted information includes the concept’s label (defined by the annotation property rdfs:label), a set of synonyms (retrieved using the annotation properties listed in Table 1), and its immediate parent concepts as defined by the rdfs:subClassOf relation.
For concepts defined using EquivalentClass axioms, the built-in verbalisation module in DeepOnto [31] is used to convert logical expressions into natural language descriptions (Table 2).
@@
To generate candidate concept pairs for alignment, GenOM adopts an embedding-based retrieval strategy. Concepts from both the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$ are first encoded into fixed-size vector representations. Each concept is represented using a combination of its label, synonyms, and its enriched definition produced in the previous step. Figure 2 shows an example of the input text fed into the embedding model. Structural information such as hierarchical relations is deliberately excluded at this stage to reduce complexity in embedding. The framework adopts the text-embedding-3-small [^3] https://platform.openai.com/docs/models/text-embedding-3-small model provided by OpenAI for embedding generation. Compared to traditional encoder-based models such as Sentence-BERT [32], this embedding model benefits from large-scale pre-training on diverse textual corpora and is better suited to capturing nuanced semantic differences expressed in longer, definition-enriched inputs, which is particularly important for distinguishing fine-grained biomedical concepts.
After generating embeddings for all target ontology concepts, we build a Hierarchical Navigable Small World (HNSW) index using the FAISS library [33]. The index is constructed in the cosine similarity space by normalising all vectors to unit length, allowing FAISS to treat cosine similarity as equivalent to inner-product distance. This HNSW structure enables efficient approximate nearest-neighbour retrieval with sublinear query complexity, substantially reducing the computational cost compared with exhaustive pairwise similarity computation.
@@
The predicted equivalence score is computed using the probability assigned to the YES token, extracted directly from the model’s output logits. This follows the probabilistic interpretation of neural network outputs widely adopted in modern confidence estimation research, where token-level softmax probabilities are treated as calibrated or approximately calibrated confidence indicators [34, 35, 36].
@@
Using token-level likelihoods as confidence estimates has been shown to correlate with model correctness in controlled-output settings, particularly when the answer space is restricted to a small set of discrete labels [35, 36]. In our case, the mapping judgement is explicitly formulated as a binary decision (YES/NO), making token probability a natural and computationally efficient confidence proxy. This probability-based scoring approach substantially reduces inference time, eliminates the need for generating full-length textual responses, and provides a straightforward decision rule while maintaining high alignment precision.

### v3

@@ Abstract
Ontology matching (OM) plays an essential role in enabling semantic interoperability and integration across heterogeneous knowledge sources, particularly in biomedical domains which contain numerous complex concepts related to diseases and pharmaceuticals. This paper introduces GenOM, a large language model (LLM)-based ontology alignment framework, which enriches semantic representations of ontology concepts via generating textual definitions, retrieving alignment candidates with an embedding model, and incorporating exact lexical matching tools to improve precision. Extensive experiments conducted on the OAEI Bio-ML track demonstrate that GenOM can often achieve competitive performance, surpassing many baselines including traditional OM systems and recent LLM-based methods. Further ablation studies confirm the effectiveness of semantic enrichment, highlighting the framework’s robustness and adaptability. Beyond the matching framework itself, this paper introduces a set of criteria for evaluating the quality of concept definitions that are generated, providing a more systematic basis for analysing LLM-generated descriptions.
@@
This paper has been accepted for publication in World Wide Web (Springer), 2026.
This version includes revisions based on peer review.
@@ 1 Introduction
In recent years, the rapid growth of domain-specific ontologies has intensified the demand for semantic interoperability and knowledge integration across heterogeneous information systems [staab2013handbook]. In particular, the biomedical domain [lambrix2009information] has witnessed the proliferation of large and complex ontologies, such as those used to model diseases, anatomy, and clinical procedures, which further amplifies this challenge. While ontologies serve as structured and machine-interpretable representations of conceptual knowledge, they are often engineered independently by different research communities or organisations. Each ontology typically reflects particular domain-specific objectives, modelling assumptions, and terminological conventions, which collectively result in substantial structural and semantic heterogeneity. This diversity, though valuable for capturing local perspectives, poses significant challenges to semantic interoperability and the effective integration of knowledge across heterogeneous information systems. Consequently, bridging conceptual discrepancies among ontologies has become a central challenge for achieving seamless knowledge exchange and reuse in data-intensive environments. Against this backdrop, this paper aims to develop a framework that leverages the reasoning and generation capabilities of large language models (LLMs) to support ontology matching across heterogeneous ontologies.
@@
Ontology matching (OM) is widely recognized as a fundamental technique for enabling semantic interoperability across heterogeneous ontologies [6104044, euzenat2013d]. Its primary objective is to uncover semantic correspondences between entities that originate from different ontologies, thereby enabling the consistent interpretation and exchange of knowledge. Through establishing formal relations such as equivalence, where two entities express the same or highly similar meanings, and subsumption, where one entity denotes a more general or more specific concept, OM serves as a cornerstone for semantic interoperability in distributed information environments.
@@
Nevertheless, achieving accurate alignments remains highly challenging due to the intrinsic heterogeneity of ontologies. This heterogeneity is manifested in several forms. First, terminological variation arises when identical concepts are described using distinct lexical labels or synonyms. Second, structural variation reflects the diversity of ontology design choices, ranging from deep hierarchical taxonomies to flat enumerative structures. Third, granularity variation occurs when similar knowledge is modelled at differing levels of abstraction. Together, these inconsistencies amplify both the cognitive and computational complexity of the matching task. The challenge becomes even more pronounced as ontologies continue to scale. For example, SNOMED CT [snomedct2024], one of the largest biomedical terminologies, comprises hundreds of thousands of concepts, which has led some studies to further decompose it into smaller, task-specific subontologies for improved tractability [delpinto2024ips, del-pinto_extracting_2022], yet purely manual alignment remains impractical at this scale, further motivating the need for scalable automated solutions.
@@
Over the past decade, a range of automated systems, such as LogMap [jimenez2011logmap] and AgreementMakerLight (AML) [faria2013agreementmakerlight], have been proposed to address this challenge. These approaches integrate lexical similarity, structural heuristics, and background knowledge resources to identify candidate mappings. While effective in well-structured and lexically rich scenarios, they often fail to capture the deeper semantics encoded in complex concept descriptions. Recent progress in LLMs has opened new possibilities for addressing this limitation [he2023exploring]. By virtue of their powerful contextual reasoning and linguistic generalisation abilities, LLMs offer an unprecedented opportunity to represent and compare ontological knowledge at a semantic level [brown2020language]. Several contemporary OM systems have begun to incorporate LLMs at different stages of the alignment pipeline, including semantic embedding and direct alignment prediction. For instance, LLM4OM [giglou2024llms4om] integrates ChatGPT with OpenAI embeddings to support pairwise concept matching, while Olala [hertling2023olala] leverages LLM-based semantic reasoning to assess concept equivalence. A more detailed review is given in Section 2. Despite these advances, current LLM-based approaches remain constrained by practical and methodological limitations. Some rely on extremely large models (e.g., the 70B-parameter LLaMA-2 employed in Olala [hertling2023olala]), which incur considerable computational costs and hinder scalability, whereas others [Norouzi2023ConversationalOA] struggle to maintain robustness when confronted with complex or large-scale alignment scenarios.
@@
To address these limitations, we develop GenOM, an ontology matching framework that leverages LLMs for semantic enrichment, candidate generation, and mapping judgement. [^1] A preliminary version of this idea, focusing on LLM-based definition generation, was presented at the OM Workshop @ ISWC 2025 [song2025genom]. GenOM begins by extracting lexical and structural information from both source and target ontologies—including labels, synonyms, parent relations, and equivalence axioms—and employs an LLM-driven description generation module to produce semantically enriched textual definitions for each concept. These enriched descriptions are then embedded into a vector space to support efficient retrieval of semantically similar candidates, while an LLM-based mapping judgement component and a lightweight exact-matching module jointly refine the alignments.
@@
In this work, we instantiate GenOM using a diverse suite of LLMs, which vary in architectural design and parameter scale. This enables a systematic analysis of how model capacity influences definition quality, semantic representation, and overall alignment performance.
@@
Contributions: This work makes the following key contributions:
  • GenOM: an LLM-enhanced ontology matching framework. We introduce GenOM, a new ontology matching framework that integrates LLM-driven semantic enrichment with embedding-based candidate generation and LLM-based mapping judgement. Unlike previous approaches that rely primarily on existing ontology annotations, GenOM exploits the conceptual knowledge implicitly stored in large language models to generate informative definitions for ontology concepts. This capability is particularly beneficial for concepts whose ontology descriptions are sparse or incomplete, for example when only a label is available. By enriching such concepts with LLM-generated definitions, GenOM provides additional semantic signals that substantially improve the ability of the downstream alignment model to distinguish correct mappings from closely related alternatives, as demonstrated in our experiments.
  • A novel evaluation methodology for assessing the quality of LLM-generated concept definitions. Since the generated definitions directly influence candidate retrieval and equivalence judgement, it is important to assess their semantic correctness and discriminative ability. We propose a dedicated evaluation scheme for concept definitions in OM tasks, combining lexical metrics, alignment-based measures, and an LLM-as-a-judge semantic assessment to quantify the correctness and discriminativeness of generated definitions.
@@ 2 Related Work
At present, OM approaches can be broadly categorised into four main types: traditional knowledge-based systems, machine learning-based systems, pre-trained language model-based systems and more recently, large language models (LLMs)-based systems. Traditional systems, such as LogMap and AML, rely primarily on lexical similarity, structural heuristics, and external resources like UMLS or WordNet [anam2015review]. LogMap [jimenez2011logmap] extracts class names and searches for matches via external lexicons while addressing logical inconsistencies by selecting alignments with higher confidence scores. AML [faria2013agreementmakerlight] employs multiple matching strategies, including exact and character-based matchers, and has shown strong performance on medical datasets. However, traditional ontology matching systems largely rely on string-based similarity measures and curated external lexicons, which have been shown to limit their coverage when semantically equivalent concepts differ substantially at the lexical level [he2022bertmap]. As a result, such approaches may fail to identify valid correspondences that are not explicitly recorded in external dictionaries or lexicons.
@@
With the rise of machine learning, several approaches have emerged that regard alignment as a classification or (embedding-based) similarity learning task [doan2004ontology, kolyvakis-etal-2018-deepalignment, nkisi2018ontology, wang2018ontology, chen2021augmenting, bento2020ontology]. For example, DeepAlignment [kolyvakis-etal-2018-deepalignment] vectorises class names and computes Euclidean distances to assess similarity, while the CNN-based system [bento2020ontology] uses character-level embeddings and hierarchical context to train binary classifiers for equivalence detection.
@@
In parallel, encoder-based pre-trained language models such as BERT [devlin-etal-2019-bert] have been explored for ontology matching, leveraging contextualised representations learned from large-scale corpora. BioSTransformers [menad2023biostransformers] adopt a Siamese architecture based on domain-specific BERT models to compute semantic similarity between biomedical concepts. BERTMap [he2022bertmap] fine-tunes a domain-specific BERT model on ontology alignment corpora mainly automatically extracted from synonyms, enabling it to capture subtle semantic differences even when lexical overlap is low. This approach has demonstrated strong results, particularly in biomedical applications. BERTSubs [DBLP:journals/www/ChenHGJDH23] takes a similar architecture as BERTMap but focuses on the subsumption relationship.
@@
More recently, the emergence of generative LLMs, including the GPT series, the Llama series, and the T5 series, has stimulated a new line of research in ontology alignment. These LLMs provide strong contextual reasoning and linguistic generalisation capabilities, making them attractive tools for capturing semantic relations that go beyond lexical similarity. A number of studies have explored prompting strategies and retrieval-augmented pipelines to exploit these capacities. For example, Norouz et al. [Norouzi2023ConversationalOA] employed GPT-4o in a prompt-driven alignment workflow, reporting strong recall, while highlighting challenges in reliably distinguishing equivalence from hierarchical relations. Yuan et al. [he2023exploring] evaluated both open- and closed-source LLMs on biomedical alignment tasks, incorporating structural cues to strengthen predictive reliability. Hybrid frameworks that combine embedding-based similarity retrieval with LLM verification have also been proposed [giglou2024llms4om, 10.14778/3712221.3712222], aiming to mitigate hallucination and improve alignment robustness. The Olala system [hertling2023olala] further integrates embedding-based filtering and post-processing modules with Llama-2, which is responsible for generating final alignment decisions. Hu et al. [hu2025matching] have explored enriching ontological information by leveraging LLMs to expand or clarify incomplete concept metadata. Their approach employs LLaMA-3.3-70B-Instruct to, for instance, transform abbreviated forms into their likely full expressions and to generate explanatory descriptions for concepts with limited textual context. The latest version of LogMap-LLM [lushnei2025large] integrates GPT-4-mini and models from the Gemini family to assist in resolving uncertain and ambiguous candidate mappings. This hybrid integration enables the system to refine difficult alignment cases that conventional lexical- or structure-based heuristics struggle to disambiguate, resulting in modest improvements over the original LogMap pipeline.
@@
Despite recent advances, existing LLM-based and LLM-augmented ontology matching approaches continue to face notable limitations. A central challenge arises from the fragmented and often sparse lexical information available within most ontologies [hertling2023olala, giglou2024llms4om, Norouzi2023ConversationalOA]. LLMs must rely heavily on labels, short synonyms, and minimally structural cues, which provide insufficient semantic context for distinguishing between closely related or hierarchically similar medical concepts. As a result, current LLM-based systems often struggle in fine-grained disambiguation scenarios [giglou2024llms4om], where subtle terminological or relational differences must be accurately interpreted.
@@ 3.1 Task Formulation
OM is the process of discovering semantic correspondences between entities that belong to different ontologies. Formally, the input of the OM task consists of two ontologies, the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$. Let $C_{\text{s}}$ and $C_{\text{t}}$ denote the sets of named concepts contained in $O_{\text{s}}$ and $O_{\text{t}}$, respectively. The goal of OM is to identify a set of mappings that link semantically related concepts across the two ontologies. Each mapping consists of a concept pair $(c_{\text{s}},c_{\text{t}})$, where $c_{\text{s}}\in C_{\text{s}}$ and $c_{\text{t}}\in C_{\text{t}}$. The output of the task is therefore a set of correspondences:
$M=\left\{(c_{\text{s}},c_{\text{t}})\mid c_{\text{s}}\in C_{\text{s}},\,c_{\text{t}}\in C_{\text{t}}\right\}$ (1)
In practical settings, the OM task can encompass various types of semantic relations, such as equivalence, subsumption, and relatedness, depending on the intended application or evaluation protocol. Accordingly, the OM task can be regarded as a process that bridges heterogeneous conceptual representations and enables interoperability of knowledge across ontologies.
@@ 3.2 The Workflow of the GenOM Framework
The overall workflow of GenOM, illustrated in Figure 1, comprises five main components. 1) Ontology data extraction collects structural and lexical information for each named concept in the source and target ontologies, including labels, synonyms, parent concepts, and equivalence axioms. This information forms the foundational representation of each concept and serves as the input for the subsequent definition generation stage. 2) An LLM-based definition generation module produces natural language definitions based on paraphrased descriptions of concepts using the extracted information. This enriches their semantic representations—particularly for entities without explicit textual descriptions [10.1145/3746252.3761533, liu_are_2026]. The generated definitions provide richer contextual signals that are later used to compute semantic similarity between concepts. 3) A candidate generation stage employs an embedding-based retrieval process to compute pairwise cosine similarities between source and target concepts using their enriched textual representations. Based on these similarity scores, the system retrieves the top-$k$ candidate mappings for each source concept, which are then passed to the next stage for equivalence judgement. 4) An LLM-based equivalence judgement module evaluates each candidate pair, determining whether the two concepts convey equivalent meanings by leveraging their generated definitions and structural context. This stage performs semantic reasoning to verify the retrieved candidate mappings and produces equivalence predictions. 5) A post-processing and result fusion stage integrates the LLM prediction scores with cosine similarities and exact lexical matching results, retaining only high-confidence correspondences. This final step combines complementary signals from semantic reasoning, embedding similarity, and lexical matching to produce the final set of ontology mappings.
@@
Both lexical and structural characteristics of each concept from the source and target ontologies are extracted and utilized to support the following alignment process. Specifically, the extracted information includes the concept’s label (defined by the annotation property rdfs:label), a set of synonyms (retrieved using the annotation properties listed in Table 1), and its immediate parent concepts as defined by the rdfs:subClassOf relation. If certain types of information were unavailable for a concept, such as synonyms or parent concepts, the corresponding field was left empty.
For concepts defined using EquivalentClass axioms, the built-in verbalisation module in DeepOnto [DBLP:journals/semweb/HeCDHAKS24] is used to convert logical expressions into natural language descriptions (Table 2). All extraction procedures were implemented using the DeepOnto framework.
@@
To generate candidate concept pairs for alignment, GenOM adopts an embedding-based retrieval strategy. Concepts from both the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$ are first encoded into fixed-size vector representations. Each concept is represented using a combination of its label, synonyms, and its enriched definition produced in the previous step. Figure 2 shows an example of the input text fed into the embedding model. Structural information such as hierarchical relations is deliberately excluded at this stage to reduce complexity in embedding. The framework adopts the text-embedding-3-small [^3] https://platform.openai.com/docs/models/text-embedding-3-small model provided by OpenAI for embedding generation. Compared to traditional encoder-based models such as Sentence-BERT [reimers2019sentence], this embedding model benefits from large-scale pre-training on diverse textual corpora and is better suited to capturing nuanced semantic differences expressed in longer, definition-enriched inputs, which is particularly important for distinguishing fine-grained biomedical concepts.
After generating embeddings for all target ontology concepts, we build a Hierarchical Navigable Small World (HNSW) index using the FAISS library [johnson2019billion]. The index is constructed in the cosine similarity space by normalising all vectors to unit length, allowing FAISS to treat cosine similarity as equivalent to inner-product distance. This HNSW structure enables efficient approximate nearest-neighbour retrieval with sublinear query complexity, substantially reducing the computational cost compared with exhaustive pairwise similarity computation.
@@
The predicted equivalence score is computed using the probability assigned to the YES token, extracted directly from the model’s output logits. This follows the probabilistic interpretation of neural network outputs widely adopted in modern confidence estimation research, where token-level softmax probabilities are treated as calibrated or approximately calibrated confidence indicators [guo2017calibration, kadavath2022language, si2022prompting].
@@
Using token-level likelihoods as confidence estimates has been shown to correlate with model correctness in controlled-output settings, particularly when the answer space is restricted to a small set of discrete labels [kadavath2022language, si2022prompting]. In our case, the mapping judgement is explicitly formulated as a binary decision (YES/NO), making token probability a natural and computationally efficient confidence proxy. This probability-based scoring approach substantially reduces inference time, eliminates the need for generating full-length textual responses, and provides a straightforward decision rule while maintaining high alignment precision.